Consumption expenditure (C) 9 000
Investment (I) 6 500
Government spending (G) 7 000
Exports (X) 1 800
Imports (Z) 2 400
Depreciation 700
Foreign payment to the rest of the
world 300
Foreign payment from the rest of the
world 250

Q.7.2.1 Calculate the value of the country’s GDE (Gross Domestic
Expenditure).
Q.7.2.2 Compute the value for the country’s GDP (Gross Domestic Product)
at market price.
(2)
Q.7.2.3 Determine the value of the country’s NNI (Net National Income) at
market price.
(3)
Q.7.2.4 If it is predicted that the GDP will increase to 22 000 in 2020,
calculate the growth rate between 2019 and 2020.
(3)
Q.7.2.5 The measure of GDP for economic growth is not always perfect.
Describe any two problems that are associated with GDP as a
measure of economic growth.
1
Expert's answer
2020-11-30T16:37:50-0500

7.2.1

gross domestic expenditure "=C+I+G=900+6500+7000=\$22500"


7.2.2

gross domestic expenditure at market price ("GDP" market price)"= C+I+G+(X-Z)"

"=9000+6500+7000+(1800-2400) = 22500-600=\$21900"


7.2.3

Net National Income ("NNI") at market price "=GDP" (market price) "+" net factor income from abroad "-" depreciation "=21900+(250-300)-7000=21900-50-7000=\$21150"


7.2.4

growth rate of "GDP= \frac{Y_{t}-Y_{t-1}}{Y_{t-1}} \times 100"

"Y_{t}(2020) =22000"

"Y_{t-1}(2019) = 21900"

growth of "GDP= \frac{22000-21900}{21900}\times 100 = 0.456\%"


7.2.5

i) The failure to account for or represent the degree of income inequality in society.

ii) The failure to indicate whether the nation's rate of growth is sustainable or not.